OpenStudy (dboyette):

Use Chebyshev’s theorem to find what percent of the values will fall between 162 and 292 for a data set with a mean of 227 and standard deviation of 13.

OpenStudy (dboyette):

i really don't understand the way the book explains this

OpenStudy (amistre64):

In elementary stats we just brushed over Cheby. But Use Chebyshev’s theorem to find what percent of the values will fall between 161 and 229 for a data set with mean of 195 and standard deviation of 17. a) The interval (161, 229) can be written as (195-2*17, 195+2*17) which is same as (Mean -k*SD, Mean +k*SD), where k =2. According to Chebyshev’s theorem, at least 1 - (1/k-squared) of the measurements will fall within (Mean -k*SD, Mean +k*SD) But 1 - (1/k-squared) = 1 - (1/2^2) = 1 – 0.25= 0.75 Thus 75 percent of the values will fall between 161 and 229 for a data set with mean of 195 and standard deviation of 17.

OpenStudy (amistre64):

This problem seems to match yours quite nicely :)

OpenStudy (amistre64):

each of your values falls at 5 sd from the mean: 227 -162 ------ 65 ; /13 = 5 Cheb thrm says: 1-(1/5^2) will work to cover the span from -5sd to 5sd 1 - 1/25 = 24/25 = .96, or 96%